OpenStudy (anonymous):

On a soccer team, 8 out of 19 players surveyed say they had two or more siblings. The league has 570 players. Which is the best prediction of the number of players in the league that have two or more siblings? a B. C. D.

OpenStudy (anonymous):

A. x/8 = 19/570 B. x/570 = 19/8 C. x/570 = 8/19 D. x/19 = 8/570

OpenStudy (anonymous):

@alyssa_nicole

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

im pretty positive thats right , im home schooled & i had this same question on an exam and made 100.

OpenStudy (anonymous):

im home schooled to this exam is hard for me

OpenStudy (anonymous):

my teachers let me use this site for my tests

OpenStudy (anonymous):

yeah luckly im pretty good in math tho, if you have any more questions just let me know.

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

okay just open a new question! (: